# A note on the improved sparse Hanson-Wright inequalities

### v1

@@ Abstract
In this paper, we establish sparse Hanson-Wright inequalities for quadratic forms of sparse $\alpha$-sub-exponential random vectors where $\alpha\in(0,2]$. When only considering the regime $0<\alpha\leq 1$, we derive a sharper sparse Hanson-Wright inequality that achieves optimality in certain special cases. These results generalize some classical Hanson-Wright inequalities without sparse structure.
33footnotetext: Department of Mathematics, Hong Kong University of Science and Technology, Clear Water Bay, Kowloon, Hong Kong, guozhengdai1997@gmail.com. 11footnotetext: Department of Mathematics, University of California Irvine, Irvine, CA 92697, USA, yiyunh4@uci.edu.22footnotetext: Department of Mathematics, Hong Kong University of Science and Technology, Clear Water Bay, Kowloon, Hong Kong, kewang@ust.hk. 44footnotetext: Department of Mathematics, University of Southern California, 3620 Vermont Avenue Los Angeles, CA 90089, USA, yizhezhu@usc.edu.
@@ 1 Introduction
Let $\xi=(\xi_{1},\cdots,\xi_{n})^{\top}$ be a random vector with independent centered entries and $A=(a_{ij})_{n\times n}$ be a fixed matrix. Exploring the concentration properties of the following quadratic random variables
$S_{A}(\xi):=\xi^{\top}A\xi=\sum_{i,j}a_{ij}\xi_{i}\xi_{j}$
is a classic topic in probability. A well-known result, proved by Hanson and Wright [7], claims that if $\xi_{i}$ are independent centered $2$-sub-exponential (sub-gaussian) random variables such that $\max_{i}\|\xi_{i}\|_{\Psi_{2}}\leq L$, then for $t\geq 0$ (the following version was provided in [15])
@@
Recently, He, Wang and Zhu [8] established sparse Hanson-Wright type inequalities in a more general setting. Specifically, let $\{\xi_{i}=\delta_{i}\cdot\zeta_{i},1\leq i\leq n\}$ be a sequence of independent random variables, where $\delta_{i}\sim\textnormal{Bernoulli}(p_{i})$ and $\zeta_{i}$ is a centered $\alpha$-sub-exponential variable independent of $\delta_{i}$, $1\leq i\leq n$. [8, Corollary 1] showed that for $\alpha>0$ and $t\geq 0$
@@
Let $\xi_{1},\cdots,\xi_{n}$ be a sequence of random variables and $f:\mathbb{R}^{n}\to\mathbb{R}$ be a measurable function. Then, $\mathbb{E}_{\xi_{i_{1}},\cdots,\xi_{i_{s}}}f(\xi_{1},\cdots,\xi_{n})$ means only taking expectation with respect to random variables $\xi_{i_{1}},\cdots,\xi_{i_{s}}$, where $\{i_{1},\cdots,i_{s}\}\subset\{1,\cdots,n\}$. Denote $\xi\sim\mathcal{W}_{s}(\alpha)$ when $\xi$ is a symmetric Weibull variable with the scale parameter $1$ and the shape parameter $\alpha$. Specifically, $-\log\mathbb{P}\{|\xi|>x\}=x^{\alpha},x\geq 0$.
@@
The rest of the paper is organized as follows. In the remaining part of Section 1, we will first present our main results. Then, we will draw comparisons between our results and existing ones to highlight the novelties and improvements. To conclude Section 1, we will pose two open questions that warrant further exploration. Section 2 is dedicated to introducing several key lemmas. These lemmas cover aspects such as the relationship between tails and moments, the decoupling inequality, the contraction principle, and concentration inequalities. They serve as fundamental tools for the subsequent theoretical derivations. In Section 4, we shall prove our main results, and in Appendix A, we offer proofs for Proposition 1.1, Corollary 1, and Lemma 2.8.
@@ 1.1 Main results
(i). When $p_{1}=\cdots=p_{n}=1$, i.e., $\delta_{i}=1$ for $1\leq i\leq n$, Theorem 1 recovers Sambale’s result [16], see (1).
(ii). When $\alpha=2$, Theorem 1 recovers the sparse Hanson-Wright inequality for sparse sub-Gaussian random vectors obtained by Zhou in [20], see (1.7).
@@
Theorems 1 and 2 assume that the matrix $A$ is symmetric. This assumption was made primarily for the convenience of presentation. Indeed, the above results can be extended to general square matrices. The only modification required is that in many places, $A$ should be replaced by $\frac{1}{2}(A+A^{\top})$.

### v2

@@ Abstract
We establish sparse Hanson–Wright inequalities for quadratic forms of sparse $\alpha$-sub-exponential random vectors with exponent parameter $\alpha\in(0,2]$. In the regime $0<\alpha\leq 1$ we derive a refined inequality that is optimal in several canonical models. These results extend the classical Hanson–Wright bound to the sparse setting. Illustrative applications include covariance matrix estimation with incomplete observations, low-rank matrix approximation under the maximum norm with sparsified sketches, and concentration inequalities for sparse $\alpha$-sub-exponential random vectors.
33footnotetext: Department of Mathematics, Hong Kong University of Science and Technology, Clear Water Bay, Kowloon, Hong Kong, guozhengdai@ust.hk. 11footnotetext: Department of Mathematics, University of California San Diego, La Jolla, CA 92093, USA, yih@ucsd.edu.22footnotetext: Department of Mathematics, Hong Kong University of Science and Technology, Clear Water Bay, Kowloon, Hong Kong, kewang@ust.hk. 44footnotetext: Department of Mathematics, University of Southern California, 3620 Vermont Avenue Los Angeles, CA 90089, USA, yizhezhu@usc.edu.
@@ 1 Introduction
The Hanson-Wright inequality quantifies the concentration of a random quadratic form around its mean. Introduced in the 1970s [14], it is now a standard device for covariance estimation, random-matrix analysis, empirical-process bounds and high-dimensional regression.
Let $\xi=(\xi_{1},\cdots,\xi_{n})^{\top}$ be a random vector with independent centered entries and $A=(a_{ij})_{n\times n}$ be a fixed matrix. The Hanson-Wright inequality is to explore the concentration properties of the following quadratic random variables
$S_{A}(\xi):=\xi^{\top}A\xi=\sum_{i,j}a_{ij}\xi_{i}\xi_{j}.$
A well-known result, proved by Hanson and Wright [14], claims that if $\xi_{i}$ are independent centered $2$-sub-exponential (sub-gaussian) random variables such that $\max_{i}\|\xi_{i}\|_{\Psi_{2}}\leq L$, then for $t\geq 0$ (the following version was provided in [30])
@@
Recently, He, Wang and Zhu [15] established sparse Hanson-Wright type inequalities in a more general setting. Specifically, let $\{\xi_{i}=\delta_{i}\cdot\zeta_{i},1\leq i\leq n\}$ be a sequence of independent random variables, where $\delta_{i}\sim\textnormal{Bernoulli}(p_{i})$ and $\zeta_{i}$ is a centered $\alpha$-sub-exponential variable independent of $\delta_{i}$, $1\leq i\leq n$. [15, Corollary 1] showed that for $\alpha>0$ and $t\geq 0$
@@
Let $\xi_{1},\cdots,\xi_{n}$ be a sequence of random variables and $f:\mathbb{R}^{n}\to\mathbb{R}$ be a measurable function. Then, $\mathbb{E}_{\xi_{i_{1}},\cdots,\xi_{i_{s}}}f(\xi_{1},\cdots,\xi_{n})$ means only taking expectation with respect to random variables $\xi_{i_{1}},\cdots,\xi_{i_{s}}$, where $\{i_{1},\cdots,i_{s}\}\subset\{1,\cdots,n\}$. Denote $\xi\sim\mathcal{W}_{s}(\alpha)$ when $\xi$ is a symmetric Weibull variable with the scale parameter $1$ and the shape parameter $\alpha$. Specifically, $\log\mathbb{P}\{|\xi|>x\}=x^{\alpha},x\geq 0$.
@@
The rest of the paper is organized as follows. In the remaining part of Section 1, we will first present our main results. Then, we will draw comparisons between our results and existing ones to highlight the novelties and improvements. To conclude Section 1, we will pose two open questions that warrant further exploration.
In Section 2, we give three applications of our concentration inequalities. The applications include: (i) covariance matrix estimation from partially observed data; (ii) low-rank matrix approximation with sparsification; and (iii) concentration phenomena for sparse $\alpha$-sub-exponential random vectors.
Section 3 is dedicated to introducing several key lemmas. These lemmas cover aspects such as the relationship between tails and moments, the decoupling inequality, the contraction principle, and concentration inequalities. They serve as fundamental tools for the subsequent theoretical derivations.
We prove our main results in Section 4 and our applications in Section 5. In the appendix, we offer proofs for Proposition 1.1, Theorem 5, Corollary 1, and Lemma 3.8.
@@ 1.1 Main results
(i). When $p_{1}=\cdots=p_{n}=1$, i.e., $\delta_{i}=1$ for $1\leq i\leq n$, Theorem 1 recovers Sambale’s result [31], see (1). (ii). When $\alpha=2$, Theorem 1 recovers the sparse Hanson-Wright inequality for sparse sub-Gaussian random vectors obtained by Zhou in [36], see (1.7).
@@
Theorems 1 and 2 assume that the matrix AA is symmetric. This assumption was made primarily for the convenience of presentation. Indeed, the above results can be extended to general square matrices. The only modification required is that in many places, AA should be replaced by $\frac{1}{2}(A+A^{\top})$.
@@ 1.2 Disscussion
2 Applications
In this section, we present several applications in covariance estimation, randomized low-rank approximation, and concentration of random vectors. Sparsity can naturally come from the fact that a high-dimensional random vector is sparse, for example, when the elements of the vector are missing at random, or when we intentionally sparsify the vector to speed up computation.
2.1 Covariance matrix estimation under missing observations
Concentration properties of the sample covariance matrix play a central role in numerous high-dimensional inference tasks, including sparse PCA, graphical model selection, discriminant analysis, and high-dimensional regression [22].
We consider the problem of estimating the covariance matrix of a random vector from i.i.d. samples, where each vector is observed with entries missing independently. In the sub-Gaussian setting, this problem was first analyzed in [26], motivated by applications in climate science, gene expression, and cosmology, where an unbiased estimator derived from the sample covariance matrix was proposed. The corresponding error bounds were later sharpened in [19]. More recently, under the $L_{4}$$L_{2}$ moment equivalence assumption, Abdalla [1] introduced an estimator that achieves both optimal dependence on the missingness parameter $p$ and the minimax optimal convergence rate under the spectral norm. However, this estimator is computationally intractable in practice.
Beyond these contributions, a number of related results on covariance estimation with missing data have appeared in the literature, primarily for sub-Gaussian random vectors; see, for example, [6, 29, 27, 20, 36, 37]. In this work, we go beyond the sub-Gaussian framework and develop concentration results for covariance estimation under heavy-tailed distributions with missing observations.
We state our assumptions as follows:
ASSUMPTION 1 (Data assumption).
Assume that the entries of $\xi\in\mathbb{R}^{m}$ are independent, centered, $\alpha$-sub-exponential random variables with $\alpha\in(0,2]$ with variance $1$. Assume
$Y=B\xi$
where $B$ is a fixed $d\times m$ matrix and $BB^{\top}=\Sigma$.
This means the data vector $Y$ satisfies a multivariate model [10, 6] as a linear transformation of $\xi^{(s)}$, which is a common assumption in high-dimensional statistics. We assume the data vector is observed partially with missing entries, defined as follows:
ASSUMPTION 2 (Missing at random).
We observe $n$ independent samples $X^{(1)},\ldots,X^{(n)}\in\mathbb{R}^{d}$, where each
$X^{(s)}=\delta^{(s)}\circ Y^{(s)},$
where
  • $\delta^{(s)}\in\{0,1\}^{d}$ are i.i.d. Bernoulli random vectors with independent entries $\delta^{(s)}_{i}\sim\mathrm{Bernoulli}(p_{i})$, $i=1,\dots,d$;
  • $Y^{(s)}\in\mathbb{R}^{m}$ are i.i.d. copies of a random vector $Y$.
The goal is to estimate $BB^{\top}=\Sigma$. A natural and unbiased estimator of $\Sigma$ is defined by
$\widehat{\Sigma}_{jk}=\begin{cases}\frac{1}{np_{j}p_{k}}\sum_{s=1}^{n}X_{j}^{(s)}X_{k}^{(s)},&\quad j\not=k,\\ \frac{1}{np_{j}}\sum_{s=1}^{n}{(X_{j}^{(s)})}^{2}&\quad j=k.\end{cases}$
This estimator is called the inverse probability weighting (IPW) estimator [20, 11, 6, 27]. When $Y$ is a sub-Gaussian random vector, the element-wise deviation of the IPW estimator has been considered in [20, 26, 27].
As soon as the dimension $d$ exceeds the sample size $n$, however, the sample covariance matrix ceases to be consistent in the operator norm. A standard remedy is to restrict attention to the “significant” coordinates, i.e. to work with a sparse sub-vector. Motivated by this practice we analyse the maximum $k$-sparse sub-matrix operator norm of the covariance matrix [22]. In particular, we would like to bound the following quantity:
$\mathrm{RIP}_{n}(k):=\sup_{\begin{subarray}{c}\theta\in\mathbb{R}^{d},\|\theta\|_{2}\leq 1\\ \|\theta\|_{0}\leq k\end{subarray}}\left|\theta^{\top}(\widehat{\Sigma}-\Sigma)\theta\right|.$
This norm is also of importance in high-dimensional linear regression due to its connections to the restricted isometry property (RIP) [7] and the restricted eigenvalue (RE) condition [4]. When $k=1$, $\mathrm{RIP}_{n}(1)$ is $\|\hat{\Sigma}-\Sigma\|_{\infty}$ and when $k=d$, $\mathrm{RIP}_{n}(1)$ is the operator norm $\|\hat{\Sigma}-\Sigma\|_{l_{2}\to l_{2}}$.
Before giving the main result of this subsection, we introduce some notations. For the vectors $p=(p_{1},\cdots,p_{d})^{\top}$ and $\theta=(\theta_{1},\cdots,\theta_{d})^{\top}$, the vector $\theta\circ\frac{1}{p}$ is defined as follows:
$\theta\circ\frac{1}{p}=(\theta_{1}/p_{1},\cdots,\theta_{d}/p_{d})^{\top}.$ (2.1)
We define the matrix $A_{\theta,p}\in\mathbb{R}^{d\times d}$ as follows:
$(A_{\theta,p})_{jk}=\begin{cases}\theta_{j}^{2}/p_{j},&j=k,\\[4.0pt] \dfrac{\theta_{j}\theta_{k}}{p_{j}p_{k}},&j\neq k.\end{cases}$
THEOREM 3.
Assume that $\max_{i}\|\xi_{i}\|_{\Psi_{\alpha}}\leq 1$. For any $t\geq 0$, with probability at least $1-2e^{-t}$,
$\mathrm{RIP}_{n}(k)$ $\lesssim_{\alpha}\frac{(t+k\log(48ed/k))^{1/2}}{\sqrt{n}}\sup_{\theta\in\Omega}\Big(\mathbb{E}\|B^{\top}\mathrm{Diag}(\delta^{(1)})A_{\theta,p}\mathrm{Diag}(\delta^{(1)})B\|_{F}^{2}\Big)^{1/2}$
$+\Big(\frac{(t+k\log(48ed/k))^{3/4}}{n^{3/4}}+\frac{(t+k\log(48ed/k))^{2/\alpha}}{n}\Big)\times$
$\quad\|B\|_{l_{2}\to l_{2}}\cdot(\mathrm{Diag}(\sqrt{p_{1}},\cdots,\sqrt{p_{d}})\cdot B\|_{F}+\|B\|_{l_{2}\to l_{2}})\cdot\sup_{\theta\in\Omega}\|\theta\circ\frac{1}{p}\|^{2}_{2},$ (2.2)
where $\Omega=\{\theta\in\mathbb{R}^{d}:\|\theta\|_{2}\leq 1,\|\theta\|_{0}\leq k\}$.
REMARK 2.1 (Comparison with [23]).
When all $p_{i}=1$, [23] studied covariance estimation for marginally sub-Weibull random vectors. Compared with their result, Theorem 3 removes the extra $\log n$ factor of the $t^{2/\alpha}$ term, and extends the analysis to the missing-observation setting; however, our assumption on the data $Y$ is stronger, whereas [23] does not require $X$ to follow a multivariate model.
REMARK $2.$
(i). By Propositions 5.1 and 5.2 below, we have
$\Big(\mathbb{E}\|B^{\top}\mathrm{Diag}(\delta^{(1)})A_{\theta,p}\mathrm{Diag}(\delta^{(1)})B\|_{F}^{2}\Big)^{1/2}$
$\lesssim$ $\|B\|_{l_{2}\to l_{2}}\cdot(\|\mathrm{Diag}(\sqrt{p_{1}},\cdots,\sqrt{p_{d}})\cdot B\|_{F}+\|B\|_{l_{2}\to l_{2}})\cdot\|\theta\circ\frac{1}{p}\|^{2}_{2}.$ (2.3)
(ii). An exact calculation of $\mathbb{E}\|B^{\top}\mathrm{Diag}(\delta^{(1)})A_{\theta,p}\mathrm{Diag}(\delta^{(1)})B\|_{F}^{2}$ is given in the Appendix E.
2.2 Low-rank approximation with sparsified sketches
Low-rank approximation is fundamental in data analysis, signal processing, and machine learning. Many large data matrices, such as user–item ratings, gene expression profiles, or kernel matrices, are approximately low-rank due to latent factors or redundancy [34]. Exploiting this structure enables efficient algorithms and reduces storage and computation in downstream tasks.
Recent work [34, 5] shows that large matrices with small spectral norm admit nearby low-rank approximations in the maximum norm, providing uniform entrywise guarantees. Building on this perspective, we propose a randomized construction using sparsified random vectors. Instead of dense Gaussian sketches, we use probes of the form $\delta\circ\xi$, where $\delta$ is a Bernoulli mask and $\xi$ is sub-Gaussian. This reduces storage and matrix–vector costs while introducing dependence on the missingness probability $p$. Our sparse Hanson–Wright inequalities yield sharp analysis, showing that sparsification preserves robust max-norm guarantees with explicit dependence on matrix coherence.
Sparsified sketches have also appeared in stochastic trace estimation [18], analyzed via sparse Hanson–Wright inequalities [36]. Our results extend this principle to low-rank approximation, demonstrating both computational savings and theoretical control.
Before stating the main result in this subsection we fix notation. For a rank-$k$ matrix $X\in\mathbb{R}^{m\times n}$ with thin SVD $X=U\Sigma V^{\top}$, the column and row-space coherences are
$\mu_{\mathrm{col}}(X)=\frac{m}{k}\max_{i\in[m]}\|U^{\top}e_{i}\|_{2}^{2},\qquad\mu_{\mathrm{row}}(X)=\frac{n}{k}\max_{j\in[n]}\|V^{\top}\tilde{e}_{j}\|_{2}^{2},$
where $e_{i}\!\in\!\mathbb{R}^{m}$ and $\tilde{e}_{j}\!\in\!\mathbb{R}^{n}$ are the standard basis vectors.
We propose a randomized algorithm to construct a low-rank approximation of a rank-$k$ matrix $X$ given in Algorithm 1. The analysis of Algorithm 1 uses the sparse Hanson-Wright inequality for a sparse sub-Gaussian random vector.
Algorithm 1 Low rank approximation with sparsified sketches
0: A rank-$k$ matrix $X\in\mathbb{R}^{m\times n}$, target rank $r$, sparsity parameter $p$.
1: Compute singular value decomposition $X=U\Sigma V^{\top}$ where $U\in\mathbb{R}^{m\times k},\Sigma\in\mathbb{R}^{k\times k}$ and $V\in\mathbb{R}^{n\times k}$. Define
$\widetilde{U}=U\Sigma^{1/2}=[\tilde{u}^{\top}_{1},\cdots,\tilde{u}^{\top}_{m}]^{\top},\quad\widetilde{V}=V\Sigma^{1/2}=[\tilde{v}^{\top}_{1},\cdots,\tilde{v}^{\top}_{n}]^{\top}.$
2: Let $Q=[Q_{1},\cdots,Q_{r}]\in\mathbb{R}^{k\times r}$ be a random matrix with independent copies of $r^{-\frac{1}{2}}\delta\circ\xi$, where $\delta\sim\mathrm{Bernoulli}(p)$ and $\xi$ is a centered sub-gaussian random variable satisfying $\|\xi\|_{\Psi_{2}}\leq 1$ and $\mathbb{E}\xi^{2}=1$. The random variable $\delta$ is independent of $\xi$.
2: Matrix $Y\in\mathbb{R}^{m\times n}$ such that $Y_{ij}=\frac{1}{p}\tilde{u}_{i}QQ^{\top}\tilde{v}_{j}^{\top}$.
THEOREM 4.
Assume that $m,n,k\in\mathbb{N}$ satisfies $k\leq\min\{m,n\}$. Let $0<p\leq 1$ and $\varepsilon>0$. Then for every matrix $X\in\mathbb{R}^{m\times n}$ of rank $k$ and any constant $r$ such that
$k\geq r\geq C_{1}\log\frac{mn}{\eta}\cdot\max\{\frac{p}{\varepsilon^{2}},\frac{1}{\varepsilon}\},$
then Algorithm 1 outputs a matrix $Y\in\mathbb{R}^{m\times n}$ of rank no more than $r$ such that with probability at least $1-\eta$,
$\|X-Y\|_{\infty}\leq\frac{k\varepsilon}{p\sqrt{mn}}\sqrt{\mu_{\mathrm{col}}\mu_{\mathrm{row}}}\cdot\|X\|_{l_{2}\to l_{2}},$ (2.4)
where $\mu_{\mathrm{col}}$ and $\mu_{\mathrm{row}}$ are the coherence of the column and row spaces of $X$.
REMARK 2.3.
Recently, Budzinskiy [5, Theorem 3.4] proved that under the condition $k\geq r\geq\frac{C_{1}\log mn}{\varepsilon^{2}}$, there exists a matrix $Y\in\mathbb{R}^{m\times n}$ of rank no more than $r$ such that
$\|X-Y\|_{\infty}\leq\frac{k\varepsilon}{\sqrt{mn}}\sqrt{\mu_{\mathrm{col}}\mu_{\mathrm{row}}}\cdot\|X\|_{l_{2}\to l_{2}}.$ (2.5)
Relative to Budzinskiy’s theorem, Theorem 4 widens the admissible range of $r$ at the cost of a looser error bound. In practice, we can trade off rank versus accuracy according to the task at hand. Our proof shows that $Y$ can be constructed based on a significantly sparser matrix $Q$, yielding substantial savings in storage and matrix–vector products compared with the dense approximation proposed in [5].
2.3 Sparse $\alpha$-sub-exponential concentration
In this subsection we establish a concentration inequality for random vectors with independent $\alpha$-sub-exponential coordinates; such bounds are widely used [30]. The derivation is standard, so the proof is relegated to the appendix.
THEOREM 5.
Assume that $\{\delta_{i},1\leq i\leq n\}$ and $\{\zeta_{i},1\leq i\leq n\}$ are two independent sequences of random variables, where $\delta_{i}\sim\textnormal{Bernoulli}(p)$ are independent and $\zeta_{i}$ are independent $\alpha$-sub-exponential random variables with mean 0 and variance $1$. Consider a random vector $\xi=(\xi_{1},\cdots,\xi_{n})$ with $\xi_{i}=\delta_{i}\cdot\zeta_{i}$. Let $A$ be an $m\times n$ fixed matrix. Then for $0<\alpha\leq 2$ and $t\geq 0$,
$\mathbb{P}\Big\{\big|\|A\xi\|_{2}-\sqrt{p}\|A\|_{F}\big|>L^{2}(\alpha)t\Big\}\leq 2\exp\Big\{-c(\alpha)\min\big\{\big(\frac{t}{\|A\|_{l_{2}\to l_{2}}}\big)^{2},\big(\frac{t}{\|A\|_{l_{2}\to l_{2}}}\big)^{\alpha}\big\}\Big\},$ (2.6)
where $L(\alpha)=\max_{i}\|\zeta_{i}\|_{\Psi_{\alpha}}$ and $c(\alpha)>0$ is a constant depending only on $\alpha$.
REMARK 2.4.
(i). In the case $0<\alpha\leq 1$ and $p=1$, Götze et al. [13, Proposition 2.1] proved that
$\mathbb{P}\Big\{\big|\|A\xi\|_{2}-\|A\|_{F}\big|>L^{2}(\alpha)t\Big\}$ (2.7)
$\leq$ $2\exp\Big\{-c(\alpha)\min\big\{\frac{t^{2}}{\|A\|^{2-\alpha}_{F}\|A\|^{\alpha}_{l_{2}\to l_{2}}},\big(\frac{t}{\|A\|_{l_{2}\to l_{2}}}\big)^{\alpha}\big\}\Big\}.$ (2.8)
Compared with their result, Theorem 5 gives an improved tail probability due to that
$\|A\|_{l_{2}\to l_{2}}^{2}\leq\|A\|^{2-\alpha}_{F}\|A\|^{\alpha}_{l_{2}\to l_{2}}.$ (2.9)
(ii). In the case $\alpha=2$ and $p=1$, Rudelson and Vershynin [30, Theorem 2.1] proved that
$\mathbb{P}\Big\{\big|\|A\xi\|_{2}-\|A\|_{F}\big|>L^{2}(\alpha)t\Big\}\leq 2\exp\Big\{-\frac{ct^{2}}{\|A\|^{2}_{l_{2}\to l_{2}}}\Big\}.$ (2.10)
Theorem 5 recovers this result in the case $\alpha=2$ and $p=1$.
REMARK 2.5.
When $\alpha=2$, we have the following inequality:
$\|\xi_{i}\|_{\Psi_{2}}\leq\|\delta_{i}\|_{\Psi_{2}}\|\zeta_{i}\|_{\Psi_{2}}\leq L\log^{-1/2}(1+p^{-1}),$
where $L=\max_{i}\|\zeta_{i}\|_{\Psi_{2}}$. This implies that $\xi_{i}$ is a sub-Gaussian random variable. Consequently, we can derive a result of the same type as in Theorem 5 from (2.10).
In particular, Rudelson and Vershynin’s result gives us:
$\mathbb{P}\left\{\left|\|A\xi\|_{2}^{2}-p\|A\|_{F}^{2}\right|>L^{2}t\right\}\leq 2\exp\left\{-\frac{c\log^{2}(1+p^{-1})pt^{2}}{\|A\|_{l_{2}\to l_{2}}^{2}}\right\}.$
Compared to this result, Theorem 5 provides a better tail probability when $p$ is close to 0. This serves as a good example to illustrate why we study sparse Hanson-Wright type inequalities.